OpenStudy (anonymous):

The engine of a sports car rotates at 5,000 revolutions per minute (rpm). Calculate the angular speed of the engine in radians per second. Use 2(pi) radians = 1 revolution

OpenStudy (saifoo.khan):

523.6 radians per second

OpenStudy (anonymous):

can you show how....

OpenStudy (anonymous):

this is new to me

OpenStudy (anonymous):

never done trig before

OpenStudy (saifoo.khan):

5000 rpm = (5000 / 60) = 83 and 1/3 rev per second. IF (and that's a big IF) each rev = 2 radians, then the angular speed is (83 and 1/3) x 2 = 166 and 2/3 rad/sec. But wait ! Don't go away. That's a bogus answer. We cannot allow a sterling example of our fine work here at Wiki.Answers to go out with the claim that 1 revolution is the same as 2 radians. It's not ! 1 revolution = 2 pi radians The true angular speed of that engine is (83 and 1/3) times (2 pi) = 523.6 radians per second (rounded)

OpenStudy (saifoo.khan):

lol, thts a problem with most college ppl. nvm, we will work thorugh it.

OpenStudy (saifoo.khan):

through*